# LAPD: Langevin-Assisted Bayesian Active Learning for Physical Discovery

### v1

@@ 4.2 Lorenz System
Consider Lorenz system of the form:
@@
with the system dynamics being governed by three parameters: the Prandtl number $\sigma$, the Rayleigh number $\rho$, and the aspect ratio $\beta$. This system is renowned for its chaotic behavior and is widely used to study complex, nonlinear dynamics in fields such as meteorology, fluid dynamics, and climate modeling. The interplay between these parameters gives rise to intricate and often unpredictable trajectories, making the Lorenz system a classic example of deterministic chaos. We simulate the system with initial conditions $\begin{bmatrix}x(0),y(0),z(0)\end{bmatrix}=\begin{bmatrix}-8,8,27\end{bmatrix}$, starting from $t=0$ and end at $t=5$ with time intervals $\Delta t=1\cdot 10^{-3}$. Here we consider the library consisted of polynomial terms up to the second order, which consists of $10$ basis: $\begin{bmatrix}1,x(t),y(t),z(t),x^{2}(t),x(t)y(t),x(t)z(t),y^{2}(t),y(t)z(t),z^{2}(t)\end{bmatrix}$. The training data is injected with 5% noise.
@@
Figure 3 shows the prediction and the UQ result with different models. The dashed vertical line at $t=3.0s$ in the plot marks the switch between in-sample prediction and out-of-sample prediction, i.e., training data are the first 3,000 time steps, and we further perform 2,500 time steps of out-of-sample prediction. Figure 3 (a) illustrates the results in which the training data is injected with 5% noise. The predictions of the trajectories are based on the results in Table 2. It is observed that for in-sample prediction, LAPD and SINDy performs much better than UQ-SINDy and that LAPD prediction aligns better with the true trajectory compared to SINDy considering the out-of-sample prediction. The 95% confidence interval does not cover the true trajectory well for out-of-sample prediction since the prediction deviation accumulates with long predictions being made by solving the initial value problem. As shown in 3 (b), when the data is noise-free, the UQ of LAPD performs well. This result indicates that in the face of rather high noise, in order to provide good prediction and UQ, the time length of the prediction shouldn’t be so long to mitigate accumulated error. One may consider collecting new data in the later time step and setting it as a new initial condition instead of still using the old initial value.
Figure 3: (a) Trajectory prediction and confidence interval with SINDY, UQ-SINDy and LAPD on Lorenz system. Training data is injected with 5% noise. The dashed vertical line marks the transition from in-sample prediction to out-of-sample prediction. The confidence interval (95%) for LAPD is illustrated as green shaded area. (b) Trajectory prediction and confidence interval with LAPD on the Lorenz system. Training data is noise-free.
Table 2: Learned coefficients with different models on Lorenz System with 5% noise.
Basis True SINDy UQ-SINDy LAPD
X Y Z X Y Z X Y Z X Y Z
1 - - - - 5.672e-1 -5.545e-1 1.855e-2 4.449e-2 -2.583e-1 - - -
x -1.000e1 2.800e1 - -1.016e1 2.714e1 - -1.096e-1 2.250e1 -3.965e-3 -9.942e0 2.797e1 -
y 1.000e1 -1.000e0 - 1.014e1 -6.566e-1 - 7.420e-1 5.070e-1 -3.649e-3 9.981e0 -9.111e-1 -
z - - -2.667e0 - - -2.655e0 2.712e-2 8.359e-3 -1.377e0 - - -2.681e0
$x^{2}$1.728e-3 -6.264e-3 1.707e-1 - - -
$xy$ 1.000e0 - - 1.001e0 -2.500e-2 -4.793e-4 8.803e-1 - - 1.001e0
$xz$1.000e0 - - -9.856e-1 - -3.124e-1 -8.601e-1 2.113e-3 - -1.007e0 -
$y^{2}$ 1.337e-2 3.945e-3 -1.076e-2 - - -
$yz$ 3.202e-1 -1.363e-2 -5.477e-4 - - -
$z^{2}$ 1.384e-3 3.450e-4 -4.770e-2 - - -
Error Bar - - 5.267e4 1.435e-3
MSE - 2.648e0 7.994e1 1.024e0
AIC - 2.940e3 1.320e4 8.386e1
@@ 4.3 Burgers’ Equation
where $u(x,t)$ represents the velocity field, and $\nu$ denotes the viscosity coefficient. In this experiment, $\nu$ takes the value of 0.1. The training data for Burgers’ Equation is generated with 256 different positions spaced equally with $x\in[-8,8]$, and 101 different time steps spaced equally with $t\in[0,10]$. The candidate library consists of 10 terms with polynomial order up to 2 with the basis being $[1,u,u_{x},u_{xx},u^{2},uu_{x},uu_{xx},u_{x}^{2},u_{x}u_{xx},u_{xx}^{2}]$. Initial condition is set to $u(0,x)=exp(-(x-3)^{2}/2)$. The training data is injected with 0.2% noise.
@@
Basis Burgers’ Equation
True SINDy UQ-SINDy LAPD(SGLD) LAPD(cyclical SGMCMC) LAPD(reSGLD)
$1$ 1.315e-3 - - -
$u$ 8.660e-2 5.908e-3 - - -
$u_{x}$2.195e-1 -1.363e-2 - - -
$u_{xx}$ 1.000e-1 5.431e-2 5.431e-2 9.161e-2 9.276e-2 9.274e-2
$u^{2}$1.960e-1 -3.109e-2 - - -
$uu_{x}$1.000e0 -2.476e-1 -9.494e-1 -9.642e-1 -9.908e-1 -9.915e-1
$uu_{xx}$ 1.921e-1 3.481e-2 - - -
$u_{x}^{2}$ 1.596e-2 - - -
$u_{x}u_{xx}$4.601e-2 - - -
$u_{xx}^{2}$ 7.756e-3 - - -
Error Bar - - 3.503e-2 2.983e-2 1.402e-4 1.218e-4
MSE - 1.448e-3 4.066e-4 9.386e-5 1.473e-5 1.421e-5
AIC - -1.690e5 -2.019e5 -2.398e5 -2.877e5 -2.886e5
Figure 4: Solution $u$ Contours of the identified Burgers’ Equation compared with ground truth (True $u$). Top: LAPD with default reSGLD as sampler Bottom: SINDy. Left Column: Contours of True $u$, Middle Column: Contours of predicted $\hat{u}$, Right Column: Contours of Error.
Figure 5: (a)Posterior distributions of identified model coefficients for Burgers’ Equation with different Langevin MCMC samplers adopted in LAPD. (b) An illustration of Error Bar (plotted in log scale) guiding threshold selection, $(0.5-0.7)$, with convection-diffusion equation.
4.4 Convection-Diffusion Equation
Convection-diffusion equation is a widely studied partial differential equation that describes the transport of a scalar quantity under the combined effects of convection and diffusion. It plays a critical role in modeling processes such as heat transfer, pollutant dispersion in fluids, and mass transfer in chemical engineering. The balance between convective transport, which moves the quantity with a flow, and diffusive transport, which acts to spread the quantity, makes it a fundamental equation for analyzing transport phenomena. The equation is expressed as:
@@ 4.4 Convection-Diffusion Equation
where $u(x,t)$ represents the transported quantity, $c$ is the convection velocity, and $D$ is the diffusion coefficient. The training data for Convection-Diffusion Equation is generated with 201 different positions spaced equally with $x\in[0,20]$, and 501 different time steps spaced equally with $t\in[0,5]$. The candidate library consists of 10 terms with polynomial order up to 2 with the basis being $[1,u,u_{x},u_{xx},u^{2},uu_{x},uu_{xx},u_{x}^{2},u_{x}u_{xx},u_{xx}^{2}]$. The training data is injected with 0.1% noise.
@@
In this experiment, we first discover the configuration of the thresholding settings. Figure 5(b) illustrates how the Error Bar value changes with different threshold settings for the Convection-Diffusion equation. The clear U-shaped curve indicates an optimal balance between model complexity and accuracy. At lower thresholds, the model includes too many terms, while at higher thresholds, essential terms are excluded, driving the model to learn poorly with higher standard deviation for coefficients. The Error Bar reaches its minimum in the range of 0.5 to 0.9, providing a data-driven approach to threshold selection in the sequential thresholding process. Subsequently, Table 4 compares the performance of SINDy, UQ-SINDy, and the proposed LAPD method in identifying the convection-diffusion equation under 0.1% noise. The LAPD framework achieves the closest alignment with the true coefficients (e.g., $u_{x}:-1.000$, $u_{xx}:0.960$), while other methods exhibit significant deviations or spurious terms (e.g., erroneous $uu_{x}$ or $u_{xx}^{2}$ coefficients). LAPD also demonstrates superior robustness, with the lowest MSE ($2.212\times 10^{-3}$), minimal posterior uncertainty ($8.613\times 10^{-4}$), and optimal AIC ($-6.157\times 10^{5}$), underscoring its efficacy in sparse, noise-resilient dynamics discovery.
@@
Figure 6: Solution $u$ Contours of the identified convection-diffusion Equation using LAPD compared with ground truth (True $u$). Top: Initial condition IC1, Bottom: Initial condition IC2. Left Column: Contours of True $u$, Middle Column: Contours of predicted $\hat{u}$, Right Column: Contours of Error.
@@
Table 4: Learned coefficients with different models on Convection-Diffusion Equation with 0.1% noise
Basis Convection-Diffusion Equation
@@
True SINDy UQ-SINDy LAPD(SGLD) LAPD(cyclical SGMCMC) LAPD(reSGLD)
$1$ 4.090e-4 - - -
$u$ 4.051e-2 - - -
$u_{x}$1.000e0 -1.265e0 -1.200e0 -9.985e-1 -1.012e0 -1.000e0
$u_{xx}$ 1.000e0 - 4.686e-1 9.514e-1 9.551e-1 9.600e-1
$u^{2}$6.343e-1 - - -
$uu_{x}$ 1.291e0 1.283e-0 - - -
$uu_{xx}$ 5.544e-1 1.460e0 - - -
$u_{x}^{2}$ 1.148e0 - - -
$u_{x}u_{xx}$8.611e-3 - - -
$u_{xx}^{2}$ 3.627e-1 9.536e-1 - - -
Error Bar - - 1.123e1 5.141e-3 1.486e-3 8.613e-4
MSE - 5.886e-2 5.782e-2 2.327e-3 2.278e-3 2.212e-3
AIC - -2.852e5 -2.870e5 -6.106e5 -6.127e5 -6.157e5
5 Experiments with Active Learning
@@
Given the limited complexity of the sample size inherent to AL settings, we employ the Metropolis adjusted Langevin algorithm (MALA) with exact gradients via LAPD, which serves as a computationally efficient single-chain alternative to reSGLD.
5.1 Lotka-Volterra System with Active Learning
@@
We evaluate the LAPD framework on the Lotka-Volterra system, following the procedure outlined in Algorithm 1. The data pool for this experiment, denoted as $\mathcal{D}$, consists of $N=10,000$ points sampled from the states generated in the 50,000 time steps. The data simulating procedure is the same as before, except for $50,000$ time steps of states are collected, with $\Delta t=5\cdot 10^{-3}$. Furthermore, 5% noise is also injected during experiments. With $\mathcal{D}$ being randomly sampled from the simulated data, it can be viewed as an irregularly sampled data set. In this synthetic experiment setting, the acquisition of the time derivative is carried out using the original regularly sampled data. However, in realistic conditions, the acquisition of the response variable (the time derivative) is computationally expensive. Here $\lambda=0,\alpha=0.5$ are adopted.
@@
Figure 7: Illustration of active learning for the LAPD method on the Lotka–Volterra system. (a) All potential sample points (light gray) and the initial 20 randomly chosen training points (dark gray) for the first round. Red points show points added over the four following iterations using the Uncertainty + Space-Filling acquisition function, whereas blue points reflect the purely Uncertainty-based selection. (b) Log-scale plot of the error bar for Random, Uncertainty, and Uncertainty + Space-Filling acquisition functions as training points are incrementally added.
@@
Training begins with an initial set of 20 randomly selected points. In each subsequent round, 10 additional points are selected via the specified acquisition functions and added to $\mathcal{M}$. Based on the acquisition function of Uncertainty+Space Filling, the threshold Tol is set to 0.8, and experiments with other acquisition functions are carried out for the same number of rounds. As depicted in Figure 7(a), combining space filling and uncertainty criteria yields a more uniform spatial distribution of the sampled points, while relying solely on uncertainty concentrates the sampling in regions of high state magnitude. Clustered data in one area have a higher chance of model overfitting and lack representativeness of the whole dataset. As shown in Figure 7(b) and Table 5, although the error bar in the round with 60 points does not differ much ($\Delta=6\cdot 10^{-3}$ on linear scale), the learned system combining Uncertainty and Space Filling yields a learned system more closely aligned with the true dynamics. Furthermore, this combined strategy illustrates the efficiency of systematic AL over random sampling. For the error bar to drop below $10^{-2}$, the random sampling strategy requires 70 data, while the combined strategy uses 30 data, reducing required measurements by around 60%. AL acquisition functions enhance the informativeness and representativeness of selected points, other than training efficiency, under the same round of training; it also improves alignment with the ground truth dynamics. As shown in Table 5, compared to randomly adding points, the Uncertainty and Uncertainty+Space-Filling strategies more accurately preserve the nonlinear structure of the ground truth, capturing both growth and predator–prey interactions without extra terms.
@@
Table 5: Comparison of true system (Lotka Volterra) and identified systems for different strategies with 60 training points
System $x_{t}$ $y_{t}$
True system $x_{t}=x-0.1xy$ $y_{t}=-1.5y+0.075xy$
Random $x_{t}=0.181+0.992x-0.099xy$ $y_{t}=-0.311-1.449y+0.073xy$
Uncertainty $x_{t}=0.902x-0.091xy$ $y_{t}=-1.339y+0.067xy$
Uncertainty+Space-Filling $x_{t}=0.998x-0.096xy$ $y_{t}=-1.423y+0.070xy$
5.2 Burgers’ Equation with Active Learning

### v2

@@ 4.2 Lorenz System
Consider the Lorenz system of the form:
@@
with the system dynamics being governed by three parameters: the Prandtl number $\sigma$, the Rayleigh number $\rho$, and the aspect ratio $\beta$. This system is renowned for its chaotic behavior and is widely used to study complex, nonlinear dynamics in fields such as meteorology, fluid dynamics, and climate modeling. The interplay between these parameters gives rise to intricate and often unpredictable trajectories, making the Lorenz system a classic example of deterministic chaos. We simulate the system with initial conditions $\begin{bmatrix}x(0),y(0),z(0)\end{bmatrix}=\begin{bmatrix}-8,8,27\end{bmatrix}$, starting from $t=0$ and end at $t=5.5$ with time intervals $\Delta t=1\times 10^{-3}$. We consider the library consisting of polynomial terms up to the second order, which consists of $10$ basis: $\begin{bmatrix}1,x(t),y(t),z(t),x^{2}(t),x(t)y(t),x(t)z(t),y^{2}(t),y(t)z(t),z^{2}(t)\end{bmatrix}$. The training data is injected with 5% additive noise.
@@
As shown in Fig. 3, when the model is trained on noisy data, the 95% prediction intervals achieve a PICP of 92% for in-sample prediction and 67% for out-of-sample prediction. The corresponding MPIW values are 0.80 for the in-sample intervals and 3.51 for the out-of-sample intervals. Relative to the mean ground-truth prediction magnitudes of 10.09 and 9.14, respectively, these interval widths indicate that the reported coverage is not achieved through trivially inflated intervals. By contrast, in panel (b), when the data are noise-free, BLADE exhibits stronger UQ performance, with the PICP of the 95% prediction intervals reaching 100%. The corresponding MPIW values are 0.37 for the in-sample intervals and 4.67 for the out-of-sample intervals. These results suggest that uncertainty quantification is substantially more reliable in the noise-free setting, whereas under high noise the prediction intervals become less well calibrated, especially for longer-horizon out-of-sample prediction. This observation also indicates that, in noisy settings, limiting the prediction horizon or periodically reinitializing from newly collected observations may help mitigate the accumulation of predictive error.
A key difference between BLADE and UQ-SINDy is the choice of sampler: BLADE uses reSGLD, whereas UQ-SINDy uses the No-U-Turn Sampler (NUTS). Since one NUTS transition and one reSGLD update are not directly comparable in terms of either statistical information content or computational structure, we use steps/s only as a measure of raw computational throughput under matched hardware, warm-up, and sampling-step budgets. To this end, we benchmark both samplers on the noise-free Lorenz dataset under identical settings. Under this benchmark, post-warm-up reSGLD sustained 25.29 steps/s, whereas NUTS sustained 1.14 steps/s. This highlights a substantial practical difference in computational throughput between the two samplers under matched implementation settings.
Figure 3: (a) Trajectory prediction and confidence interval with SINDY, UQ-SINDy, and BLADE on the Lorenz system. Training data is injected with 5% noise. The dashed vertical line marks the transition from in-sample prediction to out-of-sample prediction. The confidence interval (95%) for BLADE is illustrated as a green shaded area. (b) Trajectory prediction and confidence interval with BLADE on the Lorenz system. Training data is noise-free.
Table 2: Learned coefficients with different models on the Lorenz System with 5% noise.
Basis True SINDy UQ-SINDy BLADE X Y Z X Y Z X Y Z X Y Z 1 - - - - 5.672e-1 -5.545e-1 1.855e-2 4.449e-2 -2.583e-1 - - - x -1.000e1 2.800e1 - -1.016e1 2.714e1 - -1.096e-1 2.250e1 -3.965e-3 -9.942e0 2.797e1 - y 1.000e1 -1.000e0 - 1.014e1 -6.566e-1 - 7.420e-1 5.070e-1 -3.649e-3 9.981e0 -9.111e-1 - z - - -2.667e0 - - -2.655e0 2.712e-2 8.359e-3 -1.377e0 - - -2.681e0 $x^{2}$ - - - - - - -1.728e-3 -6.264e-3 1.707e-1 - - - $xy$ - - 1.000e0 - - 1.001e0 -2.500e-2 -4.793e-4 8.803e-1 - - 1.001e0 $xz$ - -1.000e0 - - -9.856e-1 - -3.124e-1 -8.601e-1 2.113e-3 - -1.007e0 - $y^{2}$ - - - - - - 1.337e-2 3.945e-3 -1.076e-2 - - - $yz$ - - - - - - 3.202e-1 -1.363e-2 -5.477e-4 - - - $z^{2}$ - - - - - - 1.384e-3 3.450e-4 -4.770e-2 - - - Total Error Bar - - 5.267e4 1.435e-3 MSE - 3.913e2 6.623e3 1.564e2 AIC - 5.374e4 7.924e4 4.549e4
@@ 4.3 Burgers’ Equation
where $u(x,t)$ represents the velocity field, and $\nu$ denotes the viscosity coefficient. In this experiment, $\nu$ takes the value of 0.1. The training data for Burgers’ equation is generated with 256 different positions spaced equally with $x\in[-8,8]$, and 101 different time steps spaced equally with $t\in[0,10]$. The candidate library consists of 10 terms with polynomial order up to 2, with the basis being $1,u,u_{x},u_{xx},u^{2},uu_{x},uu_{xx},u_{x}^{2},u_{x}u_{xx},u_{xx}^{2}]$. The training data is injected with 0.2% noise. For Burgers’ equation, the initial condition is set to $u(0,x)=exp(-(x-3)^{2}/2)$.
@@
Basis Burgers’ equation True SINDy UQ-SINDy BLADE(SGLD) BLADE(cyclical SGMCMC) BLADE(reSGLD) $1$ - - 1.315e-3 - - - $u$ - 8.660e-2 5.908e-3 - - - $u_{x}$ - -2.195e-1 -1.363e-2 - - - $u_{xx}$ 1.000e-1 5.431e-2 5.431e-2 9.161e-2 9.276e-2 9.274e-2 $u^{2}$ - -1.960e-1 -3.109e-2 - - - $uu_{x}$ -1.000e0 -2.476e-1 -9.494e-1 -9.642e-1 -9.908e-1 -9.915e-1 $uu_{xx}$ - 1.921e-1 3.481e-2 - - - $u_{x}^{2}$ - - 1.596e-2 - - - $u_{x}u_{xx}$ - - -4.601e-2 - - - $u_{xx}^{2}$ - - 7.756e-3 - - - Total Error Bar - - 3.503e-2 2.983e-2 1.402e-4 1.218e-4 MSE - 1.933e-3 3.490e-5 3.619e-5 2.827e-5 1.435e-5 AIC - -1.15e5 -2.653e5 -2.644e5 -2.877e5 -2.886e5
Figure 4: Solution $u$ Contours of the identified Burgers’ equation compared with ground truth (True $u$). Top: BLADE with default reSGLD as sampler Bottom: SINDy. Left Column: Contours of True $u$, Middle Column: Contours of predicted $\hat{u}$, Right Column: Contours of Error.
Figure 5: (a) Posterior distributions of identified model coefficients for Burgers’ equation with different Langevin MCMC samplers adopted in BLADE. Here we denote the predicted coefficient for $u\frac{\partial u}{\partial x}$ as $\hat{\mu}$. (b) An illustration of Total Error Bar (plotted in log scale) guiding threshold selection, (0.50.7)(0.5-0.7), with convection–diffusion equation.
4.4 Convection–Diffusion Equation
The convection–diffusion equation is a widely studied partial differential equation that describes the transport of a scalar quantity under the combined effects of convection and diffusion. It plays a critical role in modeling processes such as heat transfer, pollutant dispersion in fluids, and mass transfer in chemical engineering. The balance between convective transport, which moves the quantity with a flow, and diffusive transport, which acts to spread the quantity, makes it a fundamental equation for analyzing transport phenomena. The equation is expressed as:
@@ 4.4 Convection–Diffusion Equation
where the term $u(x,t)$ is the transported quantity, and the parameters $c$ and $D$ correspond to the convection velocity and the diffusion coefficient. The training data for convection–diffusion equation is generated with 201 different positions spaced equally with $x\in[0,20]$, and 501 different time steps spaced equally with $t\in[0,5]$. The candidate library consists of 10 terms with polynomial order up to 2 with the basis being $1,u,u_{x},u_{xx},u^{2},uu_{x},uu_{xx},u_{x}^{2},u_{x}u_{xx},u_{xx}^{2}]$. The training data is injected with 0.1% noise.
@@
In this experiment, we first discover the configuration of the thresholding settings. Figure 5(b) illustrates how the Total Error Bar value changes with different threshold settings for the convection–diffusion equation. The clear U-shaped curve indicates an optimal balance between model complexity and accuracy. At lower thresholds, the model includes too many terms, while at higher thresholds, essential terms are excluded, driving the model to learn poorly with a higher standard deviation for coefficients. The Total Error Bar reaches its minimum in the range of 0.5 to 0.9, providing a data-driven approach to threshold selection in the sequential thresholding process. Subsequently, Table 4 compares the performance of SINDy, UQ-SINDy, and the proposed BLADE method in identifying the convection–diffusion equation under 0.1% noise. The BLADE framework achieves the closest alignment with the true coefficients (e.g., $u_{x}:-1.000$, $u_{xx}:0.960$), while other methods exhibit significant deviations or spurious terms (e.g., erroneous $uu_{x}$ or $u_{xx}^{2}$ coefficients). BLADE also demonstrates superior robustness, with the lowest MSE ($1.929\times 10^{-5}$), minimal total error bar ($8.613\times 10^{-4}$), and optimal AIC ($1.093\times 10^{6}$), underscoring its efficacy in sparse, noise-resilient dynamics discovery. Figure 6 compares the ground-truth and predicted solutions under different initial conditions, along with the corresponding error maps, showing that the predictions closely match the ground truth.
@@
Figure 6: Solution $u$ Contours of the identified convection–diffusion equation using BLADE compared with ground truth (True $u$). Top: Initial condition IC1, Bottom: Initial condition IC2. Left Column: Contours of True $u$, Middle Column: Contours of predicted $\hat{u}$, Right Column: Contours of Error.
@@
Table 4: Learned coefficients with different models on convection–diffusion equation with 0.1% noise
@@
Basis convection–diffusion equation True SINDy UQ-SINDy BLADE (SGLD) BLADE (cyclical SGMCMC) BLADE (reSGLD) $1$ - - 4.090e-4 - - - $u$ - - 4.051e-2 - - - $u_{x}$ -1.000e0 -1.265e0 -1.200e0 -9.985e-1 -1.012e0 -1.000e0 $u_{xx}$ 1.000e0 - 4.686e-1 9.514e-1 9.551e-1 9.600e-1 $u^{2}$ - - -6.343e-1 - - - $uu_{x}$ - 1.291e0 1.283e-0 - - - $uu_{xx}$ - 5.544e-1 1.460e0 - - - $u_{x}^{2}$ - - 1.148e0 - - - $u_{x}u_{xx}$ - - -8.611e-3 - - - $u_{xx}^{2}$ - 3.627e-1 9.536e-1 - - - Total Error Bar - - 1.123e1 5.141e-3 1.486e-3 8.613e-4 MSE - 3.522e-4 2.053e-5 1.944e-5 1.942e-5 1.929e-5 AIC - -8.006e5 -1.087e6 -1.089e6 - 1.092e6 -1.093e6
5 Data-Efficient System Discovery with Active Learning
@@
Having demonstrated the robustness of BLADE with sufficient data in Section 4, we now evaluate its performance in data-scarce scenarios with AL. Given the limited complexity of the sample size inherent to AL settings, we employ the Metropolis adjusted Langevin algorithm (MALA) with exact gradients via BLADE, which serves as a computationally efficient single-chain alternative to reSGLD [chewi2021optimal, dwivedi2019log].
5.1 Lotka-Volterra System with AL
@@
We evaluate the BLADE framework with AL on the Lotka-Volterra system, following the procedure outlined in Algorithm 1. The pool of potential design points for this experiment, denoted $\mathcal{D}$, consists of $N=10,000$ points randomly sampled from the states generated in the 50,000 time steps. The data simulation procedure is the same as in Section 4.1, except that $50,000$ time steps of states are collected, with $\Delta t=5\cdot 10^{-4}$. Furthermore, 5% additive noise is injected during experiments. With $\mathcal{D}$ being randomly sampled from the simulated data, it can be viewed as an irregularly sampled data set. In this synthetic experiment setting, the acquisition of the time derivative is carried out using the original regularly sampled data. However, in realistic conditions, the acquisition of the response variable (the time derivative) is computationally expensive due to the irregularity of the available data. Here $\lambda=0,\alpha=0.5$ are adopted.
@@
Figure 7: Illustration of active learning for the BLADE method on the Lotka–Volterra system. (a) All potential sample points (light gray) and the initial 20 randomly chosen training points (dark gray) for the first round. Red points show points added over the four following iterations using the Uncertainty + Space-Filling acquisition function, whereas blue points reflect the purely Uncertainty-based selection. (b) Log-scale plot of the Total Error Bar for Random, Uncertainty, and Uncertainty + Space-Filling acquisition functions as training points are incrementally added.
@@
Training begins with an initial set of 20 randomly selected points. In each subsequent round, 10 additional points are selected via the specified acquisition functions and added to $\mathcal{M}$. Based on the acquisition function of Uncertainty+Space Filling, the tolerance of convergence Tol is set to 0.8, and experiments with other acquisition functions are carried out for the same number of rounds. As depicted in Figure 7(a), combining space filling and uncertainty criteria yields a more uniform spatial distribution of the sampled points, while relying solely on uncertainty concentrates the sampling in regions of high state magnitude. Clustered data in one area have a higher chance of model overfitting and lack representativeness of the whole dataset. As shown in Figure 7(b) and Table 5, although the Total Error Bar in the round with 60 points does not differ much ($\Delta=6\cdot 10^{-3}$ on linear scale), the learned system combining Uncertainty and Space Filling yields a learned system more closely aligned with the true dynamics. Furthermore, this combined strategy illustrates the efficiency of systematic AL over random sampling. For the Total Error Bar to drop below $10^{-2}$, the random sampling strategy requires 70 data, while the combined strategy uses 30 data, reducing required measurements by approximately 60%. AL acquisition functions enhance the informativeness and representativeness of selected points, other than training efficiency, under the same round of training; it also improves alignment with the ground truth dynamics. As shown in Table 5, compared to randomly adding points, the Uncertainty and Uncertainty+Space-Filling strategies more accurately preserve the nonlinear structure of the ground truth, capturing both growth and predator–prey interactions without extra terms.
@@
Table 5: Comparison of true system (Lotka-Volterra) and identified systems for different strategies with 60 training points. Our method (Uncertainty+Space-Filling) is shown bolded.
Methods $x_{t}$ $y_{t}$ True system $x_{t}=x-0.1xy$ $y_{t}=-1.5y+0.075xy$ Random $x_{t}=0.181+0.992x-0.099xy$ $y_{t}=-0.311-1.449y+0.073xy$ Uncertainty $x_{t}=0.902x-0.091xy$ $y_{t}=-1.339y+0.067xy$ Uncertainty+Space-Filling $x_{t}=0.998x-0.096xy$ $y_{t}=-1.423y+0.070xy$
5.2 Burgers’ Equation with AL